he area of a rectangular wall of a barn is 9696 square feet. Its length is 44 feet longer than twice its width. Find the length and width of the wall of the barn.

Answer by jorel1380(3719)

Let w be the width of the wall. Then its length is 2w+44. So:
w(2w+44)=9696
2w²+44w-9696=0
w²+22w-4848=0
Using the quadratic formula, we get a positive root of 59.4911341943 ft. for the width, which gives us 162.982268389 ft. as its length. ☺☺☺☺